OpenStudy (anonymous):

Ill draw the triangle Choose the best answer choice. In Triangle ABC, the measure of angle A is 45 and AC is 10. Find AB to the nearest tenth Answer A. 10 B. 7.1 C. 14.1 D. 20

OpenStudy (anonymous):



OpenStudy (anonymous):

@apoorvk

OpenStudy (anonymous):

angle A = angle B = 45 hence its isoceles triangle AC = CB = 10 units AB = sqrt(100+100) AB = sqrt(200) So its 14.1